OpenStudy (anonymous):

Find an equation relating to x and y that give you (x,y) which is equidistant from (2,7) and (3,-4). The Wording is a bit off, but the values are exact.

OpenStudy (anonymous):

\[P(2,7)\]\[Q(x,y)\] \[R(3,-4)\] \[PQ^2,and, RQ^2\]

OpenStudy (anonymous):

find

OpenStudy (anonymous):

\[(2-x)^2+(y-7)^2=PQ^2\] \[(3-x)^2+(-4-y)^2=QR^2\] \[PQ^2=RQ^2\]

OpenStudy (anonymous):

(y+4)^2

OpenStudy (anonymous):



OpenStudy (anonymous):

what is wrong with \[(x-2)^2+(y-7)^2=(x-3)^2+(y+4)^2\]?

OpenStudy (anonymous):

nothing, visual and alternate method.